a rectangle has an area of 54 square feet. if the length is 3 feet more than the width, find the dimension of the rectangle.
Answer by checkley79(3341)

L=W+3
AREA=LW
54=(W+3)W
54=W^2+3W
W^2+3W-54=0
(W-6)(W+9)=0
W-6=0
W=6 ANS.
L=6+3
L=9 ANS.
pROOF:
54=6*9
54=54